# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 2.1.2 Undirected graphs (UGs)
An undirected graph is a graph consisting of undirected edges. Suppose a random vector $X=(X_{v}:v\in\mathrm{V})$ is indexed by vertices of the undirected graph $\mathcal{G}=(\mathrm{V},\mathrm{E})$. An undirected graph $\mathcal{G}$ defines a probabilistic model for the distribution of $X$ through a set of semantics known as the Markov properties of UGs. In the literature, there are four notions of such semantics: factorization, as well as, global, local and pairwise Markov properties, which are not equivalent in general; see Lauritzen (1996, §3.2.1). The exact choice is not essential for our approach. Yet, for simplicity, in Section˜2.2 we will assume regularity conditions so that these four notions are equivalent.
@@ 2.1.3 Directed acyclic graphs (DAGs)
A directed acyclic graph is a PDAG with only directed edges. Depending on the semantics chosen, a DAG can be interpreted either as a probabilistic model or a causal model. For the purpose of this paper, we only consider DAGs as causal models since distinct DAGs (over the same set of vertices) may encode the same probabilistic model — these DAGs are called Markov equivalent — and we require the map from graph to model to be injective. For probabilistic modeling, we instead use a special type of graphs to represent the Markov equivalence classes of DAGs, which is described next.
@@ 2.1.4 Completed partially directed acyclic graphs (CPDAGs)
A CPDAG, also known as an essential graph, is a partially directed acyclic graph (PDAG) that represents a set of DAGs that are Markov equivalent. By a classical result (Verma and Pearl, 1990; Andersson, Madigan and Perlman, 1997), two DAGs are Markov equivalent if they share (1) the same skeletons (e.g., edges ignoring the directions) and (2) the same set of v-structures (i.e., unshielded colliders). We have a v-structure at a vertex $v_{2}$ if there are two edges, e.g., $v_{1}\rightarrow v_{2}$ and $v_{2}\leftarrow v_{3}$, pointing to $v_{2}$, but there is no edge between $v_{1}$ and $v_{3}$. Given a class of Markov equivalent DAGs, the CPDAG $\mathcal{G}$ representing the class can be obtained as follows. First, let $\mathcal{G}$ have the same skeleton as any (and hence every) DAG in the class. Then, for each edge in the skeleton, orient the edge accordingly if the corresponding edge points to the same direction in every DAG in the equivalence class; otherwise, let the edge be undirected. In the rest of this paper, we only consider a CPDAG $\mathcal{G}$ as the probabilistic model associated with any (and consequently every) DAG in the corresponding Markov equivalent class. It is worth mentioning that CPDAGs, or a larger type of graphs introduced below, can also be used to represent causal models. Not every PDAG is a valid CPDAG; see Section˜B.1 for details.
@@ 2.1.5 Maximally oriented partially directed acyclic graphs (MPDAGs)
A CPDAG $v_{1}-v_{2}-v_{3}$ represents three DAGs, namely $v_{1}\rightarrow v_{2}\rightarrow v_{3}$, $v_{1}\leftarrow v_{2}\leftarrow v_{3}$ and $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which are equivalent probabilistic models but distinct causal models. In the causal context, suppose we also have the background knowledge that $v_{2}$ temporally precedes $v_{3}$. Then, we can narrow down to $v_{1}\rightarrow v_{2}\rightarrow v_{3}$ or $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which together can be represented by a PDAG $v_{1}-v_{2}\rightarrow v_{3}$. PDAGs like this are MPDAGs, which represent the conjunction of several distinct causal models that are observationally Markov equivalent. They generalize CPDAGs by incorporating additional background (or experimental) knowledge on the orientation of certain edges (Hauser and Bühlmann, 2012; Perkovic, 2020; Hauser and Bühlmann, 2015). In the context of causal modeling, CPDAGs are special cases of MPDAGs (without any additional background knowledge). Therefore, in the rest of this paper, we will only focus on the probabilistic models represented by CPDAGs and the causal models represented by MPDAGs. Not every PDAG is valid MPDAG; see Section˜B.2 for structural constraints that an MPDAG must obey.
@@ 2.2 Graphs as statistical models
For the four types of graphs we introduced, we describe the models they represent. To ensure precision in deriving the corresponding model-oriented distances we do so rigorously, but the reader may set aside the technical details on first reading.
@@ 2.2.1 Graphs as probabilistic models
UGs and CPDAGs represent probabilistic models for a random vector $(X_{v}:v\in\mathrm{V})$. Suppose $X$ takes value in a product measurable space $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$. Let $\mu=\prod_{v\in V}\mu_{v}$ be a product measure over $\mathcal{X}$. Let $P$ be the law of $X$. We use the nonparametric model
@@
For UGs, as explained in Section˜2.1.2, we pose additional regularity condition on the ground set such that four Markov properties all lead to the same model definition. Let the ground set be
@@
For a CPDAG $\mathcal{G}$, it represents the probabilistic model consisting of densities that factorize according any (and hence every) DAG $\mathcal{D}$ represented by $\mathcal{G}$ (i.e., $\mathcal{D}$ is in the Markov equivalence class represented by $\mathcal{G}$), given by
@@ 2.2.2 Graphs as causal models
We follow Zhao (2025) in defining causal models associated with DAGs and MPDAGs. Consider a causal system that generates observations of a random vector $X=(X_{v}:v\in\mathrm{V})$, which takes value in a product measurable space $\mathcal{X}\equiv\mathcal{X}_{\mathrm{V}}$. Here we use notation $\mathcal{X}_{I}:=\prod_{v\in\mathrm{I}}\mathcal{X}_{v}$ for every $\mathrm{I}\subseteq\mathrm{V}$. A causal model is a set of joint distributions of counterfactual variables as well as factual (i.e., observed) variables, both of which are part of the potential outcome schedule:
$X(\cdot):=\left(X_{v}(x_{I}):v\in V,I\subseteq V,x_{I}\in\mathcal{X}_{I}\right),$
where $X_{v}(x_{I})$ is the potential outcome of variable $v$ under the intervention that imposes value $x_{I}$ on $X_{I}$. In the special case when $I=\emptyset$, the margin $(X_{v}(\emptyset):v\in V)$ consists of all the factual variables, whose law is called the observed distribution. The potential outcome schedule is a stochastic process that takes value in $\mathcal{X}^{\prod_{I\subseteq V}\mathcal{X}_{I}}$, which can be equipped with a suitable $\sigma$-algebra to form a measurable space. We use $P$ to denote the law of $X(\cdot)$ (i.e., a probability measure over the measurable space). Similar to the definition of $\mathcal{M}^{\mathrm{o}}$, we consider the ground causal model to be
$\mathcal{M}^{\mathrm{c}}:=\big\{P:P\left(X(x_{I},x_{I^{\prime}})=X(x_{I})\mid X_{I^{\prime}}(x_{I})=x_{I^{\prime}}\right)=1,\\ \text{ for every disjoint $I,I^{\prime}\subseteq V$ and every $x\in\mathcal{X}$}\big\},$ (3)
where $\mathrm{c}$’ stands for causal. In words, $\mathcal{M}^{\mathrm{c}}$ is the set of distributions of $X(\cdot)$ that maintain consistency through recursive substitution (see, e.g., Shpitser, Richardson and Robins, 2022, §41.2): in a world that sets $X_{I}$ to be $x_{I}$, if we observe $X_{I^{\prime}}$ to be $x_{I}^{\prime}$, then the observed $X$ equals $X(x_{I},x_{I^{\prime}})$. Any causal model that we discuss next will be a subset of $\mathcal{M}^{\mathrm{c}}$.
@@ A DAG $\mathcal{D}$ represents the causal model
An MPDAG $\mathcal{G}$ represents the set of Markov equivalent DAGs that also obey any extra orientation encoded by $\mathcal{G}$. A DAG $\mathcal{D}$ is represented by $\mathcal{G}$ if $\mathcal{D}$ can be obtained by orienting the undirected edges in $\mathcal{G}$ without introducing any directed cycles or v-structures. The causal model represented by MPDAG $\mathcal{G}$ is simply defined in terms of a disjunction:
@@ 3 A general recipe for model-oriented distance using partially ordered sets
Let $\mathfrak{G}$ be a finite collection of graphs over a common set of vertices $\mathrm{V}$. Let $\mathcal{M}$ be an associated model map (i.e., semantics) that assigns a model to each graph in the class. Given $(\mathfrak{G},\mathcal{M})$, our goal is to define an $\mathcal{M}$-oriented distance $d(\mathcal{G}_{s},\mathcal{G}_{t})$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. For example, for $\mathrm{V}=\{1,2,3,4,5\}$, we can consider defining distances for $(\mathfrak{G}_{1},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ and $(\mathfrak{G}_{2},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ with
[…]
@@ 3.1 Model-oriented distance
By construction, neighborhood is a symmetric relation. Intuitively, $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if it represents the simplest refinement or coarsening of $\mathcal{G}_{1}$’s model, with no graph lying strictly between them in the model hierarchy. The model-oriented distance between two graphs counts the minimal number of such basic refinement or coarsening operations to transform one model into the other, which we define as follows.
@@
The significance of the symmetry and anti-symmetry properties is straightforward. We illustrate the importance of triangle inequality from both statistical and algorithmic perspectives. From a statistical perspective, when proving consistency or convergence of an estimator, one often introduces an intermediate “ideal" object (e.g., the population version of the estimator) to bridge the gap between the data and the truth. The triangle inequality ensures that the distance between an estimated graph and the true graph can be controlled via the intermediate graph. From an algorithmic perspective, the triangle inequality helps in enumerating graphs within a small distance of a reference graph. For example, when the distance takes integer values, one can use the triangle inequality to enumerate all graphs within distance two of a reference graph by first finding all graphs at distance one, and then enumerating all graphs that are distance one from those. Without the triangle inequality, this decomposition is no longer possible, making enumeration much more expensive.
@@
For every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, there exists a finite sequence of graphs $\mathcal{G}_{s}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}=\mathcal{G}_{t}$ with $k\geq 1$ such that every $\mathcal{G}_{i},\mathcal{G}_{i+1}$ are neighbors.
@@
Let the model-oriented distance be denoted by $d$. Under Condition˜1, function $d$ can be viewed as the shortest-path distance, which is a metric, defined on an undirected graph with vertex set $\mathfrak{G}$ and edges incident to neighbors. Condition˜2 ensures that the graph is connected so $d$ is finite. For the reverse direction, if Condition˜1 fails then there exists $\mathcal{G}_{s}\neq\mathcal{G}_{t}$ with $d(\mathcal{G}_{s},\mathcal{G}_{t})=0$. If Condition˜2 fails, then there exist $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathcal{G}$ with $d(\mathcal{G}_{1},\mathcal{G}_{2})$ undefined. ∎
For natural choices of $(\mathfrak{G},\mathcal{M})$, we argue that Conditions˜1 and 2 are easily satisfied. In fact, in Section˜A.1, we show that all the four models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$ discussed in this paper satisfy Condition˜1. Typically, Condition˜2 can be ensured by the checking either of the two following stronger condition.
@@
When it exists, $\mathcal{G}_{\hat{0}}$ is the unique least element in $\mathfrak{G}$ and is usually taken to be the empty graph, which represents the smallest model. This condition has the following dual.
@@
Defining distance through the minimal number of basic operations to transform one object to the other is common in other combinatorial settings. In the context of rankings or permutations, the Kendall’s tau distance (Kendall, 1938) counts the smallest number of adjacent swaps needed to align two rankings. In clustering, the variation of information (Meila, 2005) counts the smallest weighted number of merges or splits needed to align two partitions. In both cases, the distance reflects how many basic, interpretable operations are required to transform one model into another. Here, we use the notion of neighbor to provide a parsimonious and model-oriented definition of what the basic operations are. In what follows, we use posets to formalize our approach.

### v2

@@ 2.2.2 Undirected graphs (UGs)
An undirected graph is a graph consisting of undirected edges. Suppose a random vector $X=(X_{v}:v\in\mathrm{V})$ corresponds to the vertices of the undirected graph $\mathcal{G}=(\mathrm{V},\mathrm{E})$. An undirected graph $\mathcal{G}$ defines a probabilistic model for the distribution of $X$ through a set of semantics known as the Markov properties of UGs. In the literature, there are four notions of such semantics: factorization, as well as, global, local and pairwise Markov properties, which are not equivalent in general; see Lauritzen (1996, §3.2.1). The exact choice is not essential for our approach. Yet, for simplicity, in Section˜2.3 we will assume regularity conditions so that these four notions are equivalent.
@@ 2.2.3 Directed acyclic graphs (DAGs)
A directed acyclic graph is a PDAG with only directed edges. Depending on the semantics chosen, a DAG can be interpreted either as a probabilistic model or a causal model. For the purpose of this paper, we only consider DAGs as causal models since distinct DAGs (over the same set of vertices) may encode the same probabilistic model — these DAGs are called Markov equivalent — and we require the map from graph to model to be injective. For probabilistic modeling, we instead use a special type of graph to represent the Markov equivalence classes of DAGs, which is described next.
@@ 2.2.4 Completed partially directed acyclic graphs (CPDAGs)
A CPDAG, also known as an essential graph, is a partially directed acyclic graph (PDAG) that represents a set of DAGs that are Markov equivalent. By a classical result (Verma and Pearl, 1990; Andersson, Madigan and Perlman, 1997), two DAGs are Markov equivalent if they share (1) the same skeletons (i.e., edges ignoring the directions) and (2) the same set of v-structures (i.e., unshielded colliders). We have a v-structure at a vertex $v_{2}$ if there are two edges, e.g., $v_{1}\rightarrow v_{2}$ and $v_{2}\leftarrow v_{3}$, pointing to $v_{2}$, but there is no edge between $v_{1}$ and $v_{3}$. Given a class of Markov equivalent DAGs, the CPDAG $\mathcal{G}$ representing the class can be obtained as follows. First, let $\mathcal{G}$ have the same skeleton as any (and hence every) DAG in the class. Then, for each edge in the skeleton, orient the edge accordingly if the corresponding edge points to the same direction in every DAG in the equivalence class; otherwise, let the edge be undirected. In the rest of this paper, we only consider a CPDAG $\mathcal{G}$ as the probabilistic model associated with any (and consequently every) DAG in the corresponding Markov equivalent class. It is worth mentioning that CPDAGs, or a larger type of graphs introduced below, can also be used to represent causal models. Not every PDAG is a valid CPDAG; see Section˜B.1 for details.
@@ 2.2.5 Maximally oriented partially directed acyclic graphs (MPDAGs)
A CPDAG $v_{1}-v_{2}-v_{3}$ represents three DAGs, namely $v_{1}\rightarrow v_{2}\rightarrow v_{3}$, $v_{1}\leftarrow v_{2}\leftarrow v_{3}$ and $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which are equivalent probabilistic models but distinct causal models. In the causal context, suppose we also have the background knowledge that $v_{2}$ temporally precedes $v_{3}$. Then, the set of DAGS can be narrowed down to $v_{1}\rightarrow v_{2}\rightarrow v_{3}$ and $v_{1}\leftarrow v_{2}\rightarrow v_{3}$, which together are represented by a PDAG $v_{1}-v_{2}\rightarrow v_{3}$. PDAGs like this are MPDAGs, which represent the conjunction of several distinct causal models that are observationally Markov equivalent. They generalize CPDAGs by incorporating additional background (or experimental) knowledge on the orientation of certain edges (Hauser and Bühlmann, 2012; Perkovic, 2020; Hauser and Bühlmann, 2015). In the context of causal modeling, CPDAGs are special cases of MPDAGs (without any additional background knowledge). Therefore, in the rest of this paper, we will only focus on the probabilistic models represented by CPDAGs and the causal models represented by MPDAGs. Not every PDAG is a valid MPDAG; see Section˜B.2 for structural constraints that an MPDAG must obey.
@@ 2.3 Graphs as statistical models
For the four graph types we introduced, we describe the models they represent. To derive the corresponding model-oriented distances precisely, we do so rigorously, though the reader may set aside these technical details on first reading.
@@ 2.3.1 Graphs as probabilistic models
UGs and CPDAGs represent probabilistic models for a random vector $(X_{v}:v\in\mathrm{V})$. Suppose $X$ takes values in a product measurable space $\mathcal{X}=\prod_{v\in\mathrm{V}}\mathcal{X}_{v}$. Let $\mu=\prod_{v\in\mathrm{V}}\mu_{v}$ be a product measure over $\mathcal{X}$. Let $P$ be the law of $X$. We use the nonparametric model
@@
For UGs, as explained in Section˜2.2.2, we pose an additional regularity condition on the ground set such that all four Markov properties lead to the same model definition. Concretely, for UGs let the ground set be
@@
A CPDAG $\mathcal{G}$ represents the probabilistic model consisting of densities that factorize according to any (and hence every) DAG $\mathcal{D}$ represented by $\mathcal{G}$ (i.e., $\mathcal{D}$ is in the Markov equivalence class represented by $\mathcal{G}$), given by
@@ 2.3.2 Graphs as causal models
We follow Zhao (2025) in defining causal models associated with DAGs and MPDAGs. Consider a causal system that generates observations of a random vector $X=(X_{v}:v\in\mathrm{V})$, which takes values in a product space $\mathcal{X}\equiv\mathcal{X}_{\mathrm{V}}$ equipped a suitable product $\sigma$-algebra. Here we use notation $\mathcal{X}_{\mathrm{I}}:=\prod_{v\in\mathrm{I}}\mathcal{X}_{v}$ for every $\mathrm{I}\subseteq\mathrm{V}$. A causal model is a set of joint distributions of counterfactual variables as well as factual (i.e., observed) variables, both of which are part of the potential outcome schedule:
$X(\cdot):=\left(X_{v}(x_{\mathrm{I}}):v\in\mathrm{V},\mathrm{I}\subseteq\mathrm{V},x_{\mathrm{I}}\in\mathcal{X}_{\mathrm{I}}\right),$
where $X_{v}(x_{\mathrm{I}})$ is the potential outcome of variable $v$ under the intervention that imposes value $x_{\mathrm{I}}$ on $X_{\mathrm{I}}$. In the special case when $\mathrm{I}=\emptyset$, the margin $(X_{v}(\emptyset):v\in\mathrm{V})$ consists of all the factual variables, whose law is called the observed distribution. The potential outcome schedule is a stochastic process that takes values in $\mathcal{X}^{\prod_{\mathrm{I}\subseteq\mathrm{V}}\mathcal{X}_{\mathrm{I}}}$, which can be equipped with a suitable $\sigma$-algebra to form a measurable space. We use $P$ to denote the law of $X(\cdot)$ (i.e., a probability measure over the measurable space). Similar to the definition of $\mathcal{M}^{\mathrm{o}}$, we consider the ground causal model to be
$\mathcal{M}^{\mathrm{c}}:=\big\{P:P\left(X(x_{\mathrm{I}},x_{\mathrm{I}^{\prime}})=X(x_{\mathrm{I}})\mid X_{\mathrm{I}^{\prime}}(x_{\mathrm{I}})=x_{\mathrm{I}^{\prime}}\right)=1,\\ \text{ for every disjoint $\mathrm{I},\mathrm{I}^{\prime}\subseteq\mathrm{V}$ and every $x\in\mathcal{X}$}\big\},$ (3)
where $\mathrm{c}$’ stands for causal. In words, $\mathcal{M}^{\mathrm{c}}$ is the set of distributions of $X(\cdot)$ that maintain consistency through recursive substitution (see, e.g., Shpitser, Richardson and Robins, 2022, §41.2): in a world that sets $X_{\mathrm{I}}$ to be $x_{\mathrm{I}}$, if we observe $X_{\mathrm{I}^{\prime}}$ to be $x_{\mathrm{I}^{\prime}}$, then the observed $X$ equals $X(x_{\mathrm{I}},x_{\mathrm{I}^{\prime}})$. Any causal model that we discuss next will be a subset of $\mathcal{M}^{\mathrm{c}}$.
@@ A DAG $\mathcal{D}$ represents the causal model
An MPDAG $\mathcal{G}$ represents the set of Markov equivalent DAGs that also obey any extra orientation encoded by $\mathcal{G}$. A DAG $\mathcal{D}$ is represented by $\mathcal{G}$ if $\mathcal{D}$ can be obtained by orienting the undirected edges in $\mathcal{G}$ without introducing any directed cycles or v-structures. The causal model represented by the MPDAG $\mathcal{G}$ is simply defined in terms of a disjunction:
@@ 3 A general recipe for model-oriented distance using partially ordered sets
Let $\mathfrak{G}$ be a collection of graphs over a finite vertex set $\mathrm{V}$. Let $\mathcal{M}$ be an associated model map (i.e., semantics) that assigns a model to each graph in the class. Given $(\mathfrak{G},\mathcal{M})$, our goal is to define a model-oriented distance $d(\mathcal{G}_{s},\mathcal{G}_{t})$ for every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$. By definition, $d$ depends on $\mathfrak{G}$ and $\mathcal{M}$. For example, for $\mathrm{V}=\{1,2,3,4,5\}$, we can define distances $d_{1},d_{2}$ respectively for $(\mathfrak{G}_{1},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$, $(\mathfrak{G}_{2},\mathcal{M}^{\mathrm{o}}_{\textsc{ug}})$ with
$\mathfrak{G}_{1}:=\{\text{all UGs over $\mathrm{V}$}\},\quad\mathfrak{G}_{2}:=\{\text{all acyclic UGs over $\mathrm{V}$}\}.$
@@ 3.1 Model-oriented distance
By construction, neighborhood is a symmetric relation. Intuitively, $\mathcal{G}_{2}$ is a neighbor of $\mathcal{G}_{1}$ if it represents a finest refinement or coarsening of $\mathcal{G}_{1}$’s model, in the sense that there is no graph lying strictly between them in the model hierarchy. The model-oriented distance between two graphs counts the minimal number of such basic refinement or coarsening operations to transform one model into the other, which we define as follows.
@@
The significance of the symmetry and anti-symmetry properties is straightforward. We underscore the importance of the triangle inequality from both statistical and algorithmic perspectives. Statistically, the triangle inequality is essential for proving properties like the consistency of an estimator . It allows one to bound the total error between an estimate and the truth by introducing an intermediate ’population-level’ proxy, ensuring that convergence to the proxy implies convergence to the truth. Algorithmically, the triangle inequality facilitates enumerating graphs. For instance, to enumerate all graphs within distance $k$ of a given graph $\mathcal{G}$, it suffices to expand the neighborhood of all graphs at distance k1k-1 to $\mathcal{G}$. Without this property, the enumeration can be computationally prohibitive.
@@
For every $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, there exists a sequence of graphs $\mathcal{G}_{s}=\mathcal{G}_{1},\dots,\mathcal{G}_{k}=\mathcal{G}_{t}$ with $k\geq 1$ such that every $\mathcal{G}_{i},\mathcal{G}_{i+1}$ are neighbors.
@@
Let $d$ denote the model-oriented distance. Under Condition˜1, function $d$ can be viewed as the shortest-path distance, which is a metric, defined on an undirected graph with vertex set $\mathfrak{G}$ and edges incident to neighbors. Condition˜2 ensures that the graph is connected so $d$ is finite. For the reverse direction, if Condition˜1 fails then there exists $\mathcal{G}_{s}\neq\mathcal{G}_{t}$ with $d(\mathcal{G}_{s},\mathcal{G}_{t})=0$. If Condition˜2 fails, then there exist $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathcal{G}$ with $d(\mathcal{G}_{1},\mathcal{G}_{2})$ undefined. ∎
For natural choices of $(\mathfrak{G},\mathcal{M})$, we argue that Conditions˜1 and 2 are easily satisfied. In fact, in Section˜A.1, we show that all the four models $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}},\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}},\mathcal{M}^{\mathrm{c}}_{\textsc{dag}},\mathcal{M}^{\mathrm{c}}_{\textsc{mpdag}}$ discussed in this paper satisfy Condition˜1. Typically, Condition˜2 can be ensured by the checking either of the two following stronger conditions.
@@
When it exists, $\mathcal{G}_{\hat{0}}$ is the unique least element in $\mathfrak{G}$ and is usually taken to be the empty graph, which represents the smallest model. Again, Condition 3 is satisfied by all four graph classes with $\mathcal{G}_{\hat{0}}$ being the empty graph. This condition also has the following dual.
@@
Defining distance through the minimal number of basic operations to transform one object to the other is common in other combinatorial settings. In the context of rankings or permutations, the Kendall’s tau distance (Kendall, 1938) counts the smallest number of adjacent swaps needed to align two rankings. In clustering, the variation of information (Meila, 2005) counts the smallest weighted number of merges or splits needed to align two partitions. In both cases, the distance reflects how many basic, interpretable operations are required to transform one model into another. Here, we use the notion of neighbor to provide a model-oriented definition of what the basic operations are. In what follows, we use posets to formalize our approach.